# Integrated Information Decomposition Unveils Major Structural Traits of In Silico and In Vitro Neuronal Networks

### v1

@@ Abstract
The properties of complex networked systems are a consequence of the interplay between the dynamics of their elements and the underlying topology, so to understand their behaviour it is crucial to gather information about details of their topological structure. However, in a actual networked systems such as the brain, recovering such details is not always a simple task. The reconstruction of connectivity in neuronal systems is usually done from the dynamical information encoded, e.g., in spike trains time series by measuring the transfer entropy between system elements, defining a significant pairwise interaction and use it to establish the effective connectivity among elements. The topological information recovered by these methods does not necessarily capture synaptic interaction, but rather the causal flow of information between elements. New theoretical frameworks in Information theory, such as integrated information decomposition ($\Phi$-ID), allows to explore the modes in which information can flow between parts of a system, opening a rich landscape of interactions between network topology, dynamics and information. Here, working with in silico and in vitro data, we use $\Phi$-ID to decompose the usual transfer entropy measure into the different modes in which information is transferred in a system, e.g, synergistic and unique information transfer and copy. More precisely, we demonstrate that unique information transfer is the most relevant measure to approximate structural topology details from network activity data, while the information copy only introduces ‘noise’ in the analysis. Although the retrieved network connectivity is still functional, it captures more details of the underlying structural topology by avoiding to take into account emergent high-order interactions and information redundancy between elements, which are important for the functional behavior, but obfuscate the detection of direct simple interactions between elements constituted by the structural network topology.
@@ I Introduction
In the present work, we use neuronal activity data from experimental recordings in neuronal cultures to examine, as a proof of concept, the capacity of $\Phi$-ID to decompose the information contained in the recordings on different atoms. This analysis is combined with data from simulated neuronal networks that mimic the experimental ones taking advantage of the knowledge of the ‘ground-truth topology’ of the in silico description. Our results show that the ‘unique’ information atom contains the most important information to reconstruct the ground-truth connectivity of an in silico neuronal network, and that the inferred connectivity can be used to extract interesting information on the topology of in vitro networks. Our analysis demonstrate the usefulness of $\Phi$-ID to understand information flow in neuronal networks in relation to their underlying circuitry.
Figure 1: Integrated information decomposition framework ($\Phi$-ID). (a) Time-delayed mutual information (TDMI) framework quantifies how the knowledge of a system’s current state $\mathbf{X}_{t}$ reduces the uncertainty (i.e., provides information) about future states $\mathbf{X}_{t+\tau}$. TDMI does not provide knowledge on information flow, which is realized through $\Phi$-ID. Dividing the system $X$ in two parts such that $X=\{x^{1},x^{2}\}$, we decompose TDMI into atoms of information that capture the different modes in which information is carried by the parts of a system, namely Unique (‘Un’, orange and yellow), Redundant (‘Red’, blue), and Synergistic (‘Syn’, red). (b) Through $\Phi$-ID, each of the $2^{4}$ atoms are contained in an ordered set represented as a redundancy lattice. Each node depicts an information atom that evolves from $\alpha$ to $\beta$, represented here with the respective $\Phi$-ID atom colors. Within the $\Phi$-ID framework, other usual information measures (Transfer Entropy, Information Storage, or Causal Density) can be decomposed into their fundamental atoms, giving the possibility to select just the relevant atoms that best suit an application of interest.
@@ II Information dynamics through integrated information decomposition
In the simplest case, $\Phi$-ID can be applied given a bipartition (division of system in two disconnected sub-systems) $\mathcal{B}$ of a system $\mathbf{X}=\left\{\mathbf{X}^{1},\mathbf{X}^{2}\right\}$ and a time delay $\tau$. Mutual information is decomposed then as
$I(\mathbf{X}_{t},\mathbf{X}_{t+\tau})=\sum_{\alpha,\beta\in\mathcal{A}\times\mathcal{A}}I_{\partial}^{\alpha\rightarrow\beta}\>.$ (1)
[…]
@@
An important advantage of using the $\Phi$-ID framework is the possibility of decomposing classical information measures such as Transfer Entropy into a set of core components that capture its fundamental atoms (Fig. 1). In this framework, Transfer Entropy can be expressed as
@@
Figure 2: Experimental and numerical data. (a) Top: Fluorescence image of a neuronal culture grown on a topographical substrate shaped as parallel bands, as sketched in the top-right cartoon. Bright spots are active neurons. Bottom: Positioning of $\simeq 1,400$ Regions of Interest (ROIs), set as a grid of $40\times 40$ square areas (color boxes) that cover the entire culture. (b) in silico construction of the neuronal culture. Neurons are placed with a higher connectivity probability along bands than across them. Colors highlight neurons belonging to one vertical band and their outgoing connections. The zoom in details network connectivity within a community and the presence of both excitatory (triangle) and inhibitory (circles) neurons. (c) Representative raster plots of activity for experiments (top) and simulations (bottom). Each blue dot shows an activation, either within a ROI (experiments) or a neuron (simulations).
@@
The decomposition of Eqs. (6)-(9) allows to identify and isolate the relevant atoms that best suit an application of interest and to dismiss the atoms that introduce ‘noise’ in that specific application.
@@ III Experimental and numerical background
We examined the potential of using $\Phi$-ID for connectivity estimation of experimental and numerical data derived from, or inspired by, in vitro neuronal activity recordings Soriano (2023); Montalà-Flaquer et al. (2022) (Fig. 2). As detailed in Sec. Data Availability Statement, experiments were prepared by growing cortical neurons, with typically 80% excitation and 20% inhibition, on substrates $6$ mm in diameter that incorporated a topographical modulation in the form of parallel bands —see Fig. 2(a)— which effectually favored strong neuronal connectivity along bands and a weaker connectivity across them Montalà-Flaquer et al. (2022). Spontaneous activity of the culture was then recorded, and the behavior of the system was analyzed through $1,400$ Regions of Interest (ROIs) that uniformly covered the substrate.
The numerical simulations were designed to mimic the experimental observations, by placing neurons on a bi-dimensional Euclidean space with a density and excitation-inhibition ratio similar to the one observed experimentally (see Fig. 2(b)). As detailed in Sec. Data Availability Statement, connectivity between neurons was then modeled according to rules that took into account the bands-like constraints of the substrate. The key advantage of the synthetic networks produced in this way is that the ground-truth connectivity is known. This allowed to compare the inferred connections with the known structural ones, so that the usefulness of each information atom for connectivity inference could be assessed. Moreover, it allowed to extract interesting features of the structural connectivity, such as community structure, that naturally emerged from the underlying anisotropy (see in Fig. 2(b) the colored groups of neurons), and allowed comparison to the same features extracted using the inferred connectivity by the different atoms. The procured networks were dynamically brought to life using the Izhikevich model, and the behavior of each individual neuron monitored.
Both experiments and simulations resulted in neuronal activity data that was qualitatively similar —see Fig. 2(c) —characterized by raster plots of spontaneous activity with a prominent bursting behavior, in which most of the ROIs or neurons activated together in a short time window, although they coexisted with smaller coactivations or noise-like sporadic events. These raster plots are the core data from which all information-theoretic analyses were extracted and effective connectivity quantified.
@@ IV Inferring neuronal network structure using integrated information decomposition
In order to use $\Phi$-ID to infer connectivity structure from neuronal activity data, we calculated time-delayed mutual information (TDMI) partial atoms following the $\Phi$-ID method Mediano et al. (2021) for each pair of neurons (simulations) or ROIs (experiments). Then, by grouping the atoms as described in Sec. II we obtain the TE between each pair. Following the method developed by Stetter et al. Stetter et al. (2012); Montalà-Flaquer et al. (2022), the raw measure of information of each atom and their combinations was normalized and thresholded for significance by following the z-score transformation
@@
$z\,_{IJ}=\frac{\text{TE}_{I\rightarrow J}-\left\langle TE_{\text{joint}}\right\rangle}{\sigma_{TE_{\text{joint}}}},$
@@
where $\text{TE}_{I\rightarrow J}$ is the raw TE score between any pair of nodes $I$ and $J$, $\left\langle TE_{\text{joint}}\right\rangle$ the average of the joint distribution of all input $X$ to $J$ and output $I$ to $Y$ (for any $X$ and $Y$), and $\sigma_{TE_{\text{joint}}}$ its standard deviation. The significance for information transfer, and therefore the existence of an effective connection $I\rightarrow J$, was established by setting a threshold $z_{th}$, so that $\forall_{I,J}[z_{IJ}<z_{th}\implies z_{IJ}=0]$. The remaining $z_{IJ}$ entries (significant connections) were set to $1$. The same process was performed for each partial atom and for the sum of atoms related to information transfer $I_{\partial}^{S\rightarrow U_{2}}+I_{\partial}^{U_{1}\rightarrow U_{2}}$.
@@
Such a threshold concept has been used in different studies to compare the changes in neuronal network organization in a variety of experimental and computational designs Ludl and Soriano (2020); Carola et al. (2021); Montalà-Flaquer et al. (2022), with the aim to describe key differences in overall network functional organization. However, as proposed by Stetter et al., Stetter et al. (2012), one could also consider the TE-inference approach as a way to predict the structural layout of a living neuronal culture (or its most relevant characteristics), which is in principle not completely possible. In such an effort for inferring as best as possible structural connectivity, it is clear that the choice of $z_{th}$ substantially affects the quality of reconstruction and the statistical properties of the inferred network.
@@
Thus, considering the above arguments, and in an attempt to predict interesting structural traits of neuronal cultures, we first considered the simulated data since its ground-truth connectivity is known, and explored all possible thresholds for each information atom, computing their corresponding Receiver Operating Characteristic (ROC) curve to evaluate the accuracy of the reconstruction. This allowed us to identify a range of adequate thresholds and to identify the most important atoms of information for the reconstruction of the core structural topology of the experimental data.
@@
For the numerical simulations, and after identifying the best atom, we applied the reconstruction with different thresholds in the range that we identified as adequate. Then, we measured important characteristics such as community structure to ascertain whether it approached the structural, ground-truth one. For identifying the communities we used the Louvain algorithm run in Python (NetworkX package) Hagberg, Swart, and S Chult (2008). In addition, for each explored threshold, we measured the degree distribution of effectively connected neurons as well as the distributions of effective connection distances. Overall, these comparisons allowed us to evaluate the impact of threshold selection on reconstruction quality.
@@ V.1 Determination of the relevant information atoms and threshold using in silico neuronal data
Figure 3(a) shows the ROC curves for the reconstruction of the ground-truth topology of the numerical data using the different atoms that constitute the transfer entropy measure. We observed that the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ —blue line in Fig. 3(a)— presents the best results for a low false positive ratio (higher significance threshold $z_{th}$). This atom dominates the TE measure, exhibiting on average higher values of information than the other atoms (not shown). The other atoms introduce ‘noise’ that worsens the results of the TE for low false positive ratios. However, for a higher false positive ratio (lower threshold), the synergistic transfer $I_{\partial}^{S\rightarrow U^{2}}$ (lime-colored line) contributes to improve the TE, which has better performance than the $I_{\partial}^{U^{1}\rightarrow U^{2}}$ alone. We demonstrate that this is the case by plotting the sum of $I_{\partial}^{U^{1}\rightarrow U^{2}}+I_{\partial}^{S\rightarrow U^{2}}$ (orange curve). We show how it matches with TE for lower thresholds (higher false positive ratios), and is just slightly below $I_{\partial}^{U^{1}\rightarrow U^{2}}$ for higher thresholds (low false positive ratios).
@@
Both $I_{\partial}^{U^{1}\rightarrow U^{2}}$ and $I_{\partial}^{S\rightarrow U^{2}}$ measure the information that is only transferred to part 2, which means information that in the future $t+\tau$ becomes unique to part 2. In $I_{\partial}^{S\rightarrow U^{2}}$, the information transferred depends on higher-order interactions between parts captured by the synergy measure. In this sense, it is not just the result of a direct interaction between neurons 1 and 2, but it captures the emergent information that arises from the pair Rosas et al. (2020a). Higher-order interactions have an impact on the causal flow of information, and therefore distance even more the functional topology from the structural one. On the other hand, the $I_{\partial}^{U^{1}\rightarrow U^{2}}$ atom captures information directly transmitted from 1 to 2, without accounting for higher-order interactions, which makes it a better approximation for capturing a direct simple interaction between both elements and, therefore, topological details more closely related to the structural aspects of the network.
@@
Figure 3: Accuracy of structural topology reconstruction in in silico neuronal networks. (a) Receiver Operating Characteristic (ROC) curves to quantify the accuracy of the reconstruction. Curves were calculated for the different information atoms that constitute the transfer entropy. Estimated networks are generated by including links with a calculated information score that exceeds an arbitrary threshold. ROC curves then contrast the fraction of true and false positives by comparing the inferred adjacency matrix with the ground-truth one. ROC curves show that, for a low false positive ratio, the transfer of unique information $I_{\partial}^{U_{1}\rightarrow U_{2}}$ provides the best estimate. For a higher false-positive ratio, the transfer entropy measure has a higher true-positive ratio. This effect is caused by the contribution of synergistic information transfer $I_{\partial}^{S\rightarrow U_{2}}$ (see how $I_{\partial}^{U_{1}\rightarrow U_{2}}+I_{\partial}^{S\rightarrow U_{2}}$ matches $TE_{1\rightarrow 2}$ for higher false positive values). (b) True- and false-positive ratios (TP and FP, respectively) as a function of the threshold $z_{\text{th}}$ for the Unique contribution $I_{\partial}^{U_{1}\rightarrow U_{2}}$, showing that there is an abrupt increase in FP below a critical threshold $z^{*}$ (vertical black dashed line). (c) Number of inferred links as a function of the threshold, where an abrupt change is also observed. (d) Difference between the TP and FP ratios, depicting a maximum at the transition threshold $z^{*}$.
Using the results for $I_{\partial}^{U^{1}\rightarrow U^{2}}$, we also plotted the values of true positive and false positive ratios as a function of the threshold —see Fig. 3(b)—. Here we observe that, below a critical threshold, the number of false positives increases dramatically with the increasing number of (falsely) identified links. Next, by plotting the difference between the ratios, we find a peak exactly at the critical threshold, as shown in the inset Figure 3(d). It is interesting to note that the difference between the true positives and false positives has a peak at a threshold value $z^{*}$ just above which a large increase in the number of identified links is observed. This provides a data-driven approach to define a minimal acceptable threshold even when the ground-truth topology is unknown, as any threshold below $z^{*}$ will imply an abrupt increase in false positive ratio that would add substantial spurious links onto the network.
@@
By focusing on the unique information transfer, we can next compare the obtained effective topology of the simulations with its ground-truth. Figure 4(a) shows the ground-truth network reconstructions for different thresholds. The color code indicates the community to which each node belongs. We observe that an increase in the threshold from $1$ to $4$ standard deviations does not qualitatively change the detected communities. In all cases, an approximate match between ground-truth and effective network is observed. However, we regard that an increase in the threshold slightly increases the number of communities, with some of the bigger ones actually splitting into two for the highest thresholds. This indicates that the threshold has an effect on spatial filtering. Indeed, higher thresholds favor connections between closer elements, as can be observed when measuring the probability distribution of information transfer grouped by Euclidean distance between elements (not shown). The effect of distance (or spatial embedding of neurons) on information transfer will be explored in detail in a future work.

### v2

@@ Abstract
The properties of complex networked systems arise from the interplay between the dynamics of their elements and the underlying topology. Thus, to understand their behaviour, it is crucial to convene as much information as possible about their topological organization. However, in a large systems such as neuronal networks, the reconstruction of such topology is usually carried out from the information encoded in the dynamics on the network, such as spike train time series, and by measuring the Transfer Entropy between system elements. The topological information recovered by these methods does not necessarily capture the connectivity layout, but rather the causal flow of information between elements. New theoretical frameworks, such as Integrated Information Decomposition ($\Phi$-ID), allow to explore the modes in which information can flow between parts of a system, opening a rich landscape of interactions between network topology, dynamics and information. Here, we apply $\Phi$-ID on in silico and in vitro data to decompose the usual Transfer Entropy measure into different modes of information transfer, namely synergistic, redundant or unique. We demonstrate that the unique information transfer is the most relevant measure to uncover structural topological details from network activity data, while redundant information only introduces residual information for this application. Although the retrieved network connectivity is still functional, it captures more details of the underlying structural topology by avoiding to take into account emergent high-order interactions and information redundancy between elements, which are important for the functional behavior, but mask the detection of direct simple interactions between elements constituted by the structural network topology.
@@
The structural topology describes how the elements of a networked system are connected with one another. In real systems, obtaining information about these structural details is not always possible, since usually one has access only to time-evolving data of the systems’ behavior. However, by using these data, it is possible to infer the causal relationship between elements and reconstruct the so-called effective topology, although in general it is not directly related to the underlying structural details of the network under study. By using the Integrated Information Decomposition framework in combination with in silico data, we show that it is possible to maximize the extraction of structural details from effective connectivity by selecting the appropriate mode of transfer information between network elements, to later apply the gained knowledge to extract key organizational features of experimental data in in vitro neuronal networks.
@@ I Introduction
In the present work, we use neuronal activity data from experimental recordings in neuronal cultures to examine, as a proof of concept, the capacity of $\Phi$-ID to decompose the information contained in the recordings on different atoms for the purpose of getting an insight into the structural topology of the recorded networks. This analysis is combined with data from simulated neuronal networks that mimic the experimental ones taking advantage of the knowledge of the ‘ground-truth topology’ of the in silico description. Our results show that the ‘unique’ information atom contains the most important information to reconstruct the ground-truth connectivity of an in silico neuronal network, and that the inferred connectivity can be used to extract interesting information on the topology of in vitro networks. Our analysis demonstrate the usefulness of $\Phi$-ID to understand information flow in neuronal networks in relation to their underlying circuitry.
Figure 1: Integrated Information Decomposition framework ($\Phi$-ID). (a) Time-delayed mutual information (TDMI) framework quantifies how the knowledge of a system’s current state $\mathbf{X}_{t}$ reduces the uncertainty (i.e., provides information) about future states $\mathbf{X}_{t+\tau}$. TDMI does not provide knowledge on information flow, which is realized through $\Phi$-ID. Dividing the system $X$ in two parts such that $X=\{x^{1},x^{2}\}$, we decompose TDMI into atoms of information that capture the different modes in which information is carried by the parts of a system, namely Unique (‘Un’, orange and yellow), Redundant (‘Red’, blue), and Synergistic (‘Syn’, red). (b) Through $\Phi$-ID, each of the $2^{4}$ atoms are contained in an ordered set represented as a redundancy lattice. Each node depicts an information atom that evolves from $\alpha$ to $\beta$, represented here with the respective $\Phi$-ID atom colors. Within the $\Phi$-ID framework, other usual information measures (e.g. Transfer Entropy, Information Storage or Causal Density Seth, Barrett, and Barnett (2011)) can be decomposed into their fundamental atoms, giving the possibility to select just the relevant atoms that best suit an application of interest.
@@ II Information dynamics through Integrated Information Decomposition
In the simplest case, $\Phi$-ID can be applied given a bipartition (division of system in two disconnected sub-systems) $\mathcal{B}$ of a system, such that $\mathbf{X}=\left\{\mathbf{X}^{1},\mathbf{X}^{2}\right\}$, and a time delay $\tau$. Mutual information is decomposed then as
$I(\mathbf{X}_{t},\mathbf{X}_{t+\tau})=\sum_{\alpha,\beta\in\mathcal{A}\times\mathcal{A}}I_{\partial}^{\alpha\rightarrow\beta}\>,$ (1)
where $\mathbf{X}_{t}$ represents the state of the system at time $t$. In our case, we apply $\Phi$-ID only to pairs of elements $\{x_{i},x_{j}\}$ (pairs of nodes in a neuronal network), computing the decomposition of the mutual information for each possible pair out of the $N$ elements of the network. Therefore, it should be noted that in our case, for each pair $\{x_{i},x_{j}\}$ there is only one possible bipartition.
Each information atom $I_{\partial}^{\alpha\rightarrow\beta}$ is described via a double redundancy function $I_{\cap}^{\alpha\rightarrow\beta}$ (see below) in a double redundancy lattice $\mathcal{A}\times\mathcal{A}$, as illustrated in Fig. 1. Here, $\alpha$ and $\beta$ represent the specific partial information decomposition information atom. e.g. redundant, unique or synergistic. The individual redundancy lattice is an ordered set $\mathcal{A}=\{\{\{\mathbf{X}^{1}\}\{\mathbf{X}^{2}\}\},\{\mathbf{X}^{1}\},\{\mathbf{X}^{2}\},\{\mathbf{X}^{1}\mathbf{X}^{2}\}\}$ as described in the Partial Information Decomposition (PID) frameworkBarrett (2015). Here, $\{\mathbf{X}^{1}\}\{\mathbf{X}^{2}\}$ is the redundant (denoted as Red) information contained in both partition state variables. $\{\mathbf{X}^{1}\}$ and $\{\mathbf{X}^{2}\}$ are the unique information in each variable, symbolized as Un1 and Un2, respectively. Finally, $\{\mathbf{X}^{1}\mathbf{X}^{2}\}$ is the information that emerges when both variables are considered together, that is, synergistic information Syn. The atoms $I_{\partial}^{\alpha\rightarrow\beta}$ are defined as a difference between the redundancy and the sum of lower atoms of the double redundancy lattice,
@@
An important advantage of using the $\Phi$-ID framework is the possibility of decomposing classical information measures such as Transfer Entropy (TE) into a set of core components that capture its fundamental atoms (Fig. 1). In this framework, Transfer Entropy can be expressed as
@@
Figure 2: Experimental and numerical data. (a) Top: Fluorescence image of a neuronal culture grown on a topographical substrate shaped as parallel bands, as sketched in the top-right cartoon. Bright spots are active neurons. Bottom: Positioning of $\simeq 1,400$ Regions of Interest (ROIs), set as a grid of $40\times 40$ square areas (color boxes) that cover the entire culture. (b) in silico construction of the neuronal culture. Neurons are placed with a higher connectivity probability along bands than across them. Colors highlight neurons belonging to one vertical band and their outgoing connections. The zoom in details network connectivity within a community and the presence of both excitatory (triangles) and inhibitory (circles) neurons. (c) Representative raster plots of activity for experiments (top) and simulations (bottom). Each blue dot shows an activation, either within a ROI (experiments) or a neuron (simulations).
@@
The decomposition of Eqs. (6)-(9) allows to identify and isolate the relevant atoms that best suit an application of interest and to neglect those that do not significantly contribute. The quality of the selected atoms can be tested by defining a criterion for the specific application. In our case, for example, for evaluating the reconstruction of network topology from simulated data, we use the rates of true positive (TP) and false positive (FP) links to later identify which atoms provided a high TP while maintaining a low FP.
@@ III Experimental and numerical background
We examined the potential of using $\Phi$-ID for connectivity estimation of experimental and numerical data derived from, or inspired by, in vitro neuronal activity recordings Soriano (2023); Montalà-Flaquer et al. (2022) (Fig. 2). As detailed in Appendix A, experiments were prepared by growing cortical neurons, with typically 80% excitation and 20% inhibition, on substrates $6$ mm in diameter that incorporated a topographical modulation in the form of parallel bands —see Fig. 2(a)— which effectually favored strong neuronal connectivity along bands and a weaker connectivity across them Montalà-Flaquer et al. (2022). Spontaneous activity of the culture was then recorded, and the behavior of the system was analyzed through $1,400$ Regions of Interest (ROIs) that uniformly covered the substrate.
The numerical simulations were designed to mimic the experimental observations, by placing neurons on a bi-dimensional Euclidean space with a density and excitation-inhibition ratio similar to the one observed experimentally (see Fig. 2(b)). As detailed in Appendix B, connectivity between neurons was then modeled according to rules that took into account the bands-like constraints of the substrate. The key advantage of the synthetic networks produced in this way is that the ground-truth connectivity is known. This allowed to compare the inferred connections with the known structural ones, so that the usefulness of each information atom for connectivity inference could be assessed. Moreover, it allowed to extract interesting features of the structural connectivity, such as community structure, that naturally emerged from the underlying anisotropy (see in Fig. 2(b) the colored groups of neurons), and allowed comparison to the same features extracted using the inferred connectivity by the different atoms. The procured networks were dynamically brought to life using the Izhikevich model, and the behavior of each individual neuron monitored.
Both experiments and simulations resulted in neuronal activity data that was qualitatively similar —see Fig. 2(c) —characterized by raster plots of spontaneous activity with a prominent bursting behavior, in which most of the ROIs or neurons activated together in a short time window, although they coexisted with smaller coactivations or noise-like sporadic events. These raster plots are the core data from which all information-theoretic analyses were extracted and effective connectivity quantified. Here, ‘effective connectivity’ specifically refers to ‘causal influence’ between pairs of nodes, in the sense that the knowledge about the present state of a node reduces uncertainty (or provides information) about the future states of the node it is interacting with Pearl and Mackenzie (2018).
@@ IV Inferring neuronal network structure using Integrated Information Decomposition
In order to use $\Phi$-ID to infer connectivity structure from neuronal activity data, we calculated time-delayed mutual information (TDMI) partial atoms following the $\Phi$-ID method Mediano et al. (2021) for each pair of neurons (simulations) or ROIs (experiments). Then, by grouping the atoms as described in Sec. II we obtain the TE between each pair. Following Montalà-Flaquer et al. Montalà-Flaquer et al. (2022) and Ludl et al. Ludl and Soriano (2020), the raw measure of information of each atom and their combinations was normalized and thresholded for significance by following the z-score normalization
@@
$z\,_{IJ}=\frac{\text{TE}_{I\rightarrow J}-\left\langle TE_{\text{joint}}\right\rangle}{\sigma_{TE_{\text{joint}}}},$ (11)
@@
where $\text{TE}_{I\rightarrow J}$ is the raw TE score between any pair of nodes $I$ and $J$, $\left\langle TE_{\text{joint}}\right\rangle$ the average of the joint distribution of all input $X$ to $J$ and output $I$ to $Y$ (for any $X$ and $Y$), and $\sigma_{TE_{\text{joint}}}$ its standard deviation. The significance for information transfer, and therefore the existence of an effective connection $I\rightarrow J$, was established by setting a threshold $z_{\text{th}}$, so that $\forall_{I,J}\;[\,z_{IJ}<z_{\text{th}}\implies z_{IJ}=0]$. The remaining $z_{IJ}$ entries (significant connections) were set to $1$. The same process was performed for each partial atom and for the sum of atoms related to information transfer $I_{\partial}^{S\rightarrow U_{2}}+I_{\partial}^{U_{1}\rightarrow U_{2}}$. The final set of significant effective connections was stored in a matrix $A=\{a_{IJ}\}$ for further analysis.
@@
Such a threshold concept has been used in different studies to compare the changes in neuronal network organization in a variety of experimental and computational designs Ludl and Soriano (2020); Carola et al. (2021); Montalà-Flaquer et al. (2022), with the aim to describe key differences in overall network functional organization. However, as proposed by Stetter et al., Stetter et al. (2012), one could also consider the TE-inference approach as a way to predict the structural layout of a living neuronal culture (or its most relevant characteristics), which is in principle not completely possible. In such an effort for inferring as best as possible structural connectivity, it is clear that the choice of $z_{\text{th}}$ substantially affects the quality of reconstruction and the statistical properties of the inferred network.
@@
Thus, considering the above arguments, and in an attempt to predict interesting structural traits of neuronal cultures, we first considered the simulated data, since its ground-truth connectivity is known, and explored all possible thresholds for each information atom, computing their corresponding Receiver Operating Characteristic (ROC) curves to evaluate the accuracy of their reconstructions. This allowed us to identify a range of adequate thresholds and to identify the most important atoms of information for the reconstruction of the core structural topology of the experimental data.
@@
For the numerical simulations, and after identifying the best atom, we applied the reconstruction with different thresholds in the range that we identified as adequate. Then, we measured important characteristics such as community structure to ascertain whether it approached the structural, ground-truth one. For identifying the communities we used the Louvain algorithm run in Python (NetworkX package) Hagberg, Swart, and S Chult (2008). In addition, for each explored threshold, we measured the distribution of number of inferred neighboring neurons as well as the distributions of euclidean distances between connected neurons. Overall, these comparisons allowed us to evaluate the impact of threshold selection on reconstruction quality.
@@ V.1 Determination of the relevant information atoms and threshold using in silico neuronal data
Figure 3(a) shows the ROC curves for the reconstruction of the ground-truth topology of the numerical data using the different atoms that constitute the Transfer Entropy measure. We observed that the unique information transfer $I_{\partial}^{U^{1}\rightarrow U^{2}}$ —blue line in Fig. 3(a)— presents the best results for a low false positive rate (higher significance threshold $z_{\text{th}}$). This atom dominates the TE measure, exhibiting on average higher values of information than the other atoms (not shown). The other atoms introduce ’residuals’ that worsen the results of the TE for low false positive rates. However, for a higher false positive rate (lower threshold), the synergistic transfer $I_{\partial}^{S\rightarrow U^{2}}$ (lime-colored line) contributes to improve the TE, which has better performance than the $I_{\partial}^{U^{1}\rightarrow U^{2}}$ alone. We demonstrate that this is the case by plotting the sum of $I_{\partial}^{U^{1}\rightarrow U^{2}}+I_{\partial}^{S\rightarrow U^{2}}$ (orange curve). We show how it matches with TE for lower thresholds (higher false positive rates), and is just slightly below $I_{\partial}^{U^{1}\rightarrow U^{2}}$ for higher thresholds (low false positive rates).
@@
Both $I_{\partial}^{U^{1}\rightarrow U^{2}}$ and $I_{\partial}^{S\rightarrow U^{2}}$ measure the information that is only transferred to part 2, which means information that in the future $t+\tau$ becomes unique to part 2. In $I_{\partial}^{S\rightarrow U^{2}}$, the information transferred depends on higher-order interactions between parts captured by the synergy measure. In this sense, it is not just the result of a direct interaction between neurons 1 and 2, but it captures the emergent information that arises from the pair Rosas et al. (2020). Higher-order interactions have an impact on the causal flow of information, and therefore depart even more the effective topology from the structural one. On the other hand, the $I_{\partial}^{U^{1}\rightarrow U^{2}}$ atom captures information directly transmitted from 1 to 2, without accounting for higher-order interactions, which makes it a better approximation for capturing a direct simple interaction between both elements and, therefore, topological details more closely related to the structural aspects of the network.
@@
Figure 3: Accuracy of structural topology reconstruction in in silico neuronal networks. (a) Receiver Operating Characteristic (ROC) curves to quantify the accuracy of the reconstruction. Curves were calculated for the different information atoms that constitute the Transfer Entropy. Estimated networks are generated by including links with a calculated information score that exceeds an arbitrary threshold. ROC curves then contrast the fraction of true and false positives by comparing the inferred adjacency matrix with the ground-truth one. ROC curves show that, for a low false positive rate, the transfer of unique information $I_{\partial}^{U_{1}\rightarrow U_{2}}$ provides the best estimate. For a higher false-positive rate, the Transfer Entropy measure has a higher true-positive rate. This effect is caused by the contribution of synergistic information transfer $I_{\partial}^{S\rightarrow U_{2}}$ (see how $I_{\partial}^{U_{1}\rightarrow U_{2}}+I_{\partial}^{S\rightarrow U_{2}}$ matches $TE_{1\rightarrow 2}$ for higher false positive values). (b) True- and false-positive rates (TP and FP, respectively) as a function of the threshold $z_{\text{th}}$ for the Unique contribution $I_{\partial}^{U_{1}\rightarrow U_{2}}$, showing that there is an abrupt increase in FP below a critical threshold $z^{*}\simeq 0.4$ (vertical black dashed line). Conceptually, this abrupt change indicates that, below $z^{*}$, the FP rate (blue curve) quickly grows with just a small increase in TP (red curve), overall providing worse reconstructions. (c) Number of inferred links as a function of the threshold, where an abrupt change is also observed. (d) Difference between the TP and FP rate, depicting a maximum at the critical threshold $z^{*}$. The shadow areas indicate the standard deviation of the values in three different realization of the in silico neuronal network. The white star indicates the critical threshold point $z^{*}$.
Using the results for $I_{\partial}^{U^{1}\rightarrow U^{2}}$, we also plotted the values of true positive and false positive rates as a function of the threshold, as shown in Fig. 3(b). Here we observe that, below a ‘critical threshold’ $z^{*}\simeq 0.4$, the number of false positives abruptly grows with just a small increase in true positives, substantially worsening reconstruction. Thresholds above $z^{*}$ smoothly reduce the amount of true positives for a very small reduction of the few remaining false positives. Indeed, just below this threshold, a large increase in the number of identified links is observed, as visible in Fig. 3(c). Additionally, by plotting the difference between the TP and FP rates, we observe a peak exactly at this $z^{*}$, as shown in Fig. 3(d). Thus, this threshold $z^{*}$ can be considered an optimal threshold in terms of the difference between TP and FP.
The point on the ROC curve of the TE reconstruction for this threshold $z^{*}$ is indicated by a white star in Fig. 3(a). It can be seen that this point coincides with the intersection of the $I_{\partial}^{U^{1}\to U^{2}}$ and the TE ROC curves, meaning that for a more restrictive threshold $z>z^{*}$ the $I_{\partial}^{U^{1}\to U^{2}}$ reconstruction outperforms the TE reconstruction. We note that the identification of $z^{*}$ by plotting the number of observed links as a function of the threshold, as in Fig. 3(c), depends only on the recorded activity dynamics and it can be therefore also applied to in vitro, experimental data. More specifically, the critical threshold $z^{*}$ appears at the right-most knee of the links-threshold curve of Fig. 3(c).
@@
By focusing on the unique information transfer, we can next compare the obtained effective topology of the simulations with its ground-truth. Figure 4(a) shows the ground-truth network reconstructions for different thresholds $z_{\text{th}}>z^{*}$. The color code indicates the community to which each node belongs. We observe that an increase in the threshold from $1$ to $4$ standard deviations does not qualitatively change the detected communities, as also revealed by a similar value of the modularity statistic $Q$. In all cases, an approximate match between ground-truth and effective network is observed. However, we regard that an increase in the threshold slightly increases the number of communities, with some of the bigger ones actually splitting into two for the highest thresholds. This gradual fragmentation of the network is expected since less connections are present overall, as captured by the global efficiency $G_{E}$, and therefore the reconstructed networks are expected to be more modular (higher $Q$). We also note that the threshold has an effect on spatial filtering, with higher thresholds favoring connections between closer elements, as observed previously by Montalà-Flaquer et al. Montalà-Flaquer et al. (2022) when measuring the probability distribution of information transfer grouped by Euclidean distance between elements. The detailed effect of distance (or spatial embedding of neurons) on information transfer will be explored in a future work.